# Trust Me, I Know the Way: Predictive Uncertainty in the Presence of Shortcut Learning

### v1

@@ 1 Introduction
We do not aspire to resolve the philosophical dilemma but argue instead that the two EU manifestations arise from different scenarios. More precisely, we postulate that conflicting hypotheses (V2) only emerge in the presence of shortcut learning (SCL): One hypothesis at most recovers the true mapping while the others necessarily point to spurious patterns. Our experiments suggest that shortcuts in the data indeed prompt such disagreement. This observation bears important insights for learning dynamics and has not, to the best of our knowledge, been studied despite ample discussion about robust generalization (Nagarajan et al., 2021; Wald et al., 2021; Richens and Everitt, 2024).
@@ 3 Predictive Uncertainty in the Presence of Shortcut Learning
In short, we observe that uncertainty estimates based on data with explicit shortcuts differ from those without. We consider classification tasks derived from MNIST (LeCun et al., 1998, pooling digits 1-3, 4-6, 7-9) and solved using deep ensembles of size 3 (Lakshminarayanan et al., 2017, see App. A.2 for details). Shortcuts of strength $s$ are introduced in $s$of samples by coloring entire images (CMNIST3) or adding $1\times 1$ colored pixels (PMNIST3) according to class777 Our code will be made available upon publication. (Fig. 1).
@@
Fig. 2 shows the distribution of uncertainty estimates for MNIST3, CMNIST3 and PMNIST3 ($s\in\{50,95,100\}$). The respective test data consist of original class-0 images not seen during training. Uncertainty under SCL manifests as near-maximal mutual information888 We normalize all uncertainty components to values in [0, 1] according to Eq. 2. (orange bars) for CMNIST3, at near-full confidence (meaning zero AU; blue bars), under strong-enough shortcuts (columns 3-4). In other words, the ensemble members produce conflicting predictions with high probability on different classes for most test images (less so when the shortcut is only partial). This behavior follows the disagreement interpretation of EU (V2 in Sec. 1). There is a similar, if weaker, pattern for the less-pronounced PMNIST3 cue. Hypotheses start to agree again later in a sort of model collapse for PMNIST3 with $s=100$ (note that estimates evolve as training progresses; see also App. A.2.2). The shortcut-free MNIST3 task produces a diffuse distribution in both uncertainty components, reflecting varying degrees of confidence and disagreement across different test images (V1). Only few samples provoke full disagreement. Recall, however, that the test images are OOD for all datasets. Using EU estimates as a signal in OOD detection would then yield quite different interpretations depending on whether SCL is at play. Taking uncertainty estimates at face value can thus be deceptive without a clear understanding of what EU via disagreement implies.
@@ 4 Conclusion
Much of recent research in UQ concerns methodological improvement. While indisputably important, this narrow focus—when it entails silent acceptance of assumptions like i.i.d. data—isolates the field from more fundamental discussions. The confusion about the appropriate representation of EU arises from such a gap. We reconcile different views on representing EU (conflict vs ignorance) by showing how shortcuts, which would not exist if the i.i.d. assumption actually held, affect uncertainty estimates in support of the conflict-type notion. With this, we hope to contribute to a comprehensive view on generalization that stands the test of real-world learning situations.
@@ Acknowledgments
This paper is supported by the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the German Federal Ministry of Education and Research.
@@ A.1.2 Random Variables
where the logarithm is typically set to base 2 in accordance with an information-theoretic bit interpretation. Entropy captures the potential information gain from observing the realization of a RV. Consequently, it is minimal for a RV whose distribution is a Dirac measure, since the outcome is all but certain a priori, and maximal for uniformly distributed RVs [Cover and Thomas, 2006].
@@ A.2.1 Experimental Details
Our code is mainly based on PyTorch [Paszke et al., 2019] and PyTorch Lightning [Falcon, 2023]. We performed all experiments on CPU and will make our code available upon publication.

### v2

@@ 1 Introduction
We do not aspire to resolve the philosophical dilemma but argue instead that the two EU manifestations might arise from different scenarios. More precisely, we postulate that conflicting hypotheses (V2) emerge in the presence of shortcut learning (SCL): Since there is only one true mapping from features to targets, one hypothesis at most—or even none—can have recovered this relationship, while the others necessarily point to spurious patterns. Our experiments suggest that shortcuts in the data indeed prompt such disagreement. This observation bears important insights for learning dynamics and has not, to the best of our knowledge, been studied despite ample discussion about robust generalization (Nagarajan et al., 2021; Wald et al., 2021; Richens and Everitt, 2024). We emphasize that we do not claim SCL is the sole cause of conflicting hypotheses (see Sec. 4); rather, our results should be taken as encouraging further research in this direction.
@@ 3 Predictive Uncertainty in the Presence of Shortcut Learning
In short, we observe that uncertainty estimates based on data with explicit shortcuts differ from those without. We consider classification tasks derived from MNIST (LeCun et al., 1998, pooling digits 1-3, 4-6, 7-9) and solved using deep ensembles of size 3 (Lakshminarayanan et al., 2017, see App. A.2 for details). Shortcuts of strength $s$ are introduced in $s$of samples by coloring entire images (CMNIST3) or adding $1\times 1$ colored pixels (PMNIST3) according to class777 Our code is available at https://github.com/lisa-wm/shortcuts_uncertainty.git. (Fig. 1).
@@
Fig. 2 shows the distribution of uncertainty estimates for MNIST3, CMNIST3 and PMNIST3 ($s\in\{50,95,100\}$). The respective test data consist of original class-0 images not seen during training. Uncertainty under SCL manifests as near-maximal mutual information888 We normalize all uncertainty components to values in [0, 1] according to Eq. 2. (orange bars) for CMNIST3, at near-full confidence (meaning zero AU; blue bars), under strong-enough shortcuts (columns 3-4). In other words, the ensemble members produce conflicting predictions with high probability on different classes for most test images (less so when the shortcut is only partial). This behavior follows the disagreement interpretation of EU (V2 in Sec. 1). There is a similar, if weaker, pattern for the less-pronounced PMNIST3 cue. Hypotheses start to agree again later in a sort of model collapse for PMNIST3 with $s=100$ (note that estimates evolve as training progresses; see also App. A.2.2). The shortcut-free MNIST3 task produces a diffuse distribution in both uncertainty components, reflecting varying degrees of confidence and disagreement across different test images (V1). Only few samples provoke full disagreement. Recall, however, that the test images are OOD for all datasets. Using EU estimates as a signal in OOD detection would then yield quite different interpretations depending on whether SCL is at play: trained on MNIST3, the evidence pointing toward OOD is rather weak, in contrast to training on data infused with a shortcut, even though both scenarios are, in fact, OOD. Taking uncertainty estimates at face value can thus be deceptive without a clear understanding of what EU via disagreement implies.
4 Discussion
The evidence presented in Sec. 3 suggests a relation between SCL and predictive uncertainty that has so far received little attention. It does not, however, imply a one-to-one correspondence between SCL and high EU. For one, situations other than SCL might prompt conflicting hypotheses: when the data are truly ambiguous or even contradictory, we must expect models to pick up different signals. That said, such poor-quality data should manifest in higher AU and thus lower EU (due to the additivity of the components and the entropy being upper-bounded), reflecting the entanglement between both sources of uncertainty in finite-data settings pointed out by Wimmer et al. (2023). On the other hand, SCL might occur without any conflict if all hypotheses come to agree on the same, shortcut-driven mapping. Our results provide initial insights in what seems to be a promising direction, and underline the importance of studying the impact of distribution shifts on phenomena usually viewed in the i.i.d. vacuum.
@@ 5 Conclusion
Much of recent research in UQ concerns methodological improvement. While indisputably important, this narrow focus—when it entails silent acceptance of assumptions like i.i.d. data—isolates the field from more fundamental discussions. The confusion about the appropriate representation of EU arises from such a gap. We take a step toward reconciling different views on representing EU (conflict vs ignorance) by showing how shortcuts, which would not exist if the i.i.d. assumption actually held, affect uncertainty estimates in support of the conflict-type notion. With this, we hope to contribute to a comprehensive view on generalization that stands the test of real-world learning situations.
@@ Acknowledgments
LW is supported by the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the German Federal Ministry of Education and Research.
@@ A.1.2 Random Variables
where the logarithm is typically set to base 2 in accordance with an information-theoretic bit interpretation. Entropy captures the potential information gain from observing the realization of RVs. Consequently, it is minimal for RVs whose distribution is a Dirac measure, since the outcome is all but certain a priori, and maximal for uniformly distributed RVs [Cover and Thomas, 2006]. Normalization of $H(A)$ by $(\log|\Omega|)^{-1}$ maps entropy values to the unit interval.
@@ A.2.1 Experimental Details
Our code is mainly based on PyTorch [Paszke et al., 2019] and PyTorch Lightning [Falcon, 2023]. We performed all experiments on CPU and make our code available at https://github.com/lisa-wm/shortcuts_uncertainty.git.
@@ A.2.2 Further Results
In-Distribution Results
Finally, we provide in Fig. 5 the results corresponding to Fig. 2 with in-distribution test data (in the sense that both partitions contain the same classes, as opposed to the OOD experiments based on predictions for the held-out 0 class). Shortcuts are present only during training. As might be expected, the disagreement between hypotheses is generally less pronounced, suggesting that the models pick up some of the non-spurious patterns. For CMNIST3 with perfect shortcut, we still see high EU values (note, though, that Fig. 5 only depicts results for the 20% most-uncertain predictions).
Figure 5: In-distribution test data. Uncertainty estimates (20% least-confident predictions) based on data with varying $s$, where train and test classes coincide but shortcuts are only present in the training data; ensemble size 3; average over 3 independent runs (error bars: $\pm$ 1 standard error).